SATURATION TRANSFER STUDIES ON DOORWAY STATES FOR CH_{2} INTERSYSTEM CROSSING

YANGSOO KIM^{a}, Chemistry Department, Brookhaven National Laboratory, Upton, NY 11973-5000; ANATOLY V. KOMISSAROV, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA 91109-8099; GREGORY E. HALL, and \underline{TREVOR J. SEARS}, Chemistry Department, Brookhaven National Laboratory, Upton, NY 11973-5000.

Double-resonance transient photobleaching kinetics of singlet and mixed singlet-triplet eigenstates of CH_{2} to characterize the closely related processes of rotational energy transfer and collision induced intersystem crossing in methylene will be reported. Qualitative features of the state-to-state rotational energy transfer matrix are revealed. Polarization dependence (parallel and perpendicular) of saturation recovery is observed. Limited by the detection sensitivity, it was not possible to identify polarization dependence of saturation transfer. Selective bleaching of individual eigenstates of mixed state (doorway) pairs shows very efficient collisional interconversion within the mixed-state pair.

aWork at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.